OpenStudy (anonymous):

Help me please r + 0.12 = 0.20 A. r= -0.08 B. r= 0.08 C. r= -0.32 D. r= 0.32

OpenStudy (nethercreep333):

Let's solve your equation step-by-step. r+0.12=0.2 Step 1: Subtract 0.12 from both sides. r+0.12−0.12=0.2−0.12 r=0.08 Answer: r=0.08